Please help me to solve this word problem:
If the average cost per unit C(x) (in dollars) to produce x units of plywood is given by C(x)=1500/(x+50, what is the unit cost for 40 units?
a) $37.50 b) $0.75 c) -$12.50 d) $16.67
Thank you.

Answer by ewatrrr(24785)

 
Hi,
If unit cost for 40 units C(x) (in dollars) to produce x units of plywood
is given by C(x)=1500/(x+50)
unit cost for 40 units is C(40) = 1500/90 = $16.67